OpenStudy (yanasidlinskiy):

A car is traveling at 33.0 m/s when the driver uses the brakes to slow down to 29.0 m/s in 2.50 seconds. How many meters will it travel during that time?

OpenStudy (sweetburger):

I think you use this kinematic equation. idk not totally sure \[distance=v_i +.5at^2\] your given time, initial velocity, and you can solve for acceleration given you have the change in velocity over a set time.

OpenStudy (sweetburger):

should work

OpenStudy (nincompoop):

we can derive it if you're not sure how the formula is figured out

OpenStudy (yanasidlinskiy):

Either way I wanna know.

OpenStudy (nincompoop):

what do you know at this point?

OpenStudy (yanasidlinskiy):

Well I know the givens. But I'm not sure which equation should I use and which would be easier. I don't know how to derive it either.

OpenStudy (nincompoop):

so what equations do you know ?

OpenStudy (yanasidlinskiy):

Well I'm using what Buger told me lol

OpenStudy (nincompoop):

smh, besides that?

OpenStudy (yanasidlinskiy):

None.

OpenStudy (nincompoop):

no clue about velocity nor acceleration?

OpenStudy (yanasidlinskiy):

Nope

OpenStudy (nincompoop):

LAUGHING OUT LOUD! for real?

OpenStudy (yanasidlinskiy):

lol yea hahha

OpenStudy (nincompoop):

try this vid https://www.youtube.com/watch?v=N6NPDoIoJWY and we can expand later.

OpenStudy (yanasidlinskiy):

KK

OpenStudy (nincompoop):

it is a short and good summary

OpenStudy (yanasidlinskiy):

ya

OpenStudy (yanasidlinskiy):

ok soo....

OpenStudy (nincompoop):

watch it and try to learn

OpenStudy (nincompoop):

do some algebra

OpenStudy (yanasidlinskiy):

Velocity Initial is 33.0m/s right?

OpenStudy (yanasidlinskiy):

I got 82.5 Is that right??

OpenStudy (nincompoop):

correct about the initial velocity

OpenStudy (yanasidlinskiy):

Wht about the answer?

OpenStudy (nincompoop):

I didn't solve it

OpenStudy (yanasidlinskiy):

Oh.. Well do it?? Lol

OpenStudy (abhisar):

Hi @YanaSidlinskiy ! How do you want me to help you?

OpenStudy (yanasidlinskiy):

Is tht answer right?

OpenStudy (abhisar):

one min

OpenStudy (yanasidlinskiy):

Ok

OpenStudy (abhisar):

Ok, the car should travel 77.5 meters

OpenStudy (yanasidlinskiy):

so 82.5 is wrong?

OpenStudy (abhisar):

Yes it is.

OpenStudy (abhisar):

There are two ways you can solve this question.

OpenStudy (yanasidlinskiy):

Will you show me how??

OpenStudy (abhisar):

\(\sf Method~I\) Use V=U+at to find the acceleration and then substitute this acceleration in any of the other equations of motion to find distance. \(\sf Method~II\) You must have noticed that the above method involves two steps and is often lengthy. So I give you a trick equation use \(\sf S= \frac{U+V}{2} \times t\) You see no info about acceleration is required here

OpenStudy (yanasidlinskiy):

I did method 1

OpenStudy (abhisar):

That's fine. Do it again. It must be a calculation error.

OpenStudy (yanasidlinskiy):

Ok well the time is obviously gonna be 2.50 IV= 33.0

OpenStudy (yanasidlinskiy):

Right?

OpenStudy (abhisar):

correct

OpenStudy (yanasidlinskiy):

I don't get what would be the U though...

OpenStudy (abhisar):

U = initial velocity, V = final velocity

OpenStudy (yanasidlinskiy):

Ohhh wow! Ok lol. So dumb. Ok and acceleration would be 29.0 because it slowed down which deaccelerated correct?

OpenStudy (abhisar):

Nopes!! Initial velocity is 33 m/s. Final velocity is 29 m/s because that's the velocity of the car after slowing down. You need to find out the acceleration using these data.

OpenStudy (yanasidlinskiy):

So?? 29m/s= 33m/s(2.50m/s)?

OpenStudy (abhisar):

How?

OpenStudy (yanasidlinskiy):

Lol idk. I'm sooooo confused rn

OpenStudy (abhisar):

You know that, Final velocity = Initial velocity + acceleration x time

OpenStudy (abhisar):

Now tell me do you think you wrote correct equation?

OpenStudy (yanasidlinskiy):

No

OpenStudy (abhisar):

Then rectify it

OpenStudy (yanasidlinskiy):

Idk what else to do to it

OpenStudy (nincompoop):

throwing in formulas do not help after all

OpenStudy (yanasidlinskiy):

Nope not at all.

OpenStudy (abhisar):

It's simple yana!! Just plug in the values dear. 29 = 33 + a x 2.5

OpenStudy (abhisar):

Getting this?

OpenStudy (yanasidlinskiy):

Ok so acceleration is only when the rate or speed increases right?

OpenStudy (abhisar):

Just be relaxed, everyone has to struggle initially. ;) Kinematics is very easy..you'll see this gradually

OpenStudy (yanasidlinskiy):

Lol stop winking. Anywho... Ok can you walk me through this problem at least

OpenStudy (nincompoop):

can we do this gradually like from the simplest of the simplest

OpenStudy (nincompoop):

I feel like you failed to understand some crucial concepts prior to learning acceleration

OpenStudy (abhisar):

"Acceleration" this term is used when speed increases and when speed decreases the term used is retardation.

OpenStudy (abhisar):

But overall both are same physical quantity.

OpenStudy (yanasidlinskiy):

Ok.

OpenStudy (nincompoop):

decelerate and accelerate not retardation

OpenStudy (yanasidlinskiy):

Ok smarty

OpenStudy (nincompoop):

I will create a post for you later

OpenStudy (yanasidlinskiy):

Thank you.

OpenStudy (abhisar):

@nincompoop , both term are used

OpenStudy (nincompoop):

all this formula memorization is rubbish

OpenStudy (yanasidlinskiy):

Ok how would I continue to solve the problem?

OpenStudy (abhisar):

Yes, find the unknown a from the above equation.

OpenStudy (yanasidlinskiy):

You really think I'm gonna read tht? Lol. And Abhi I don't know how!!!!!

OpenStudy (abhisar):

It's a simple linear equation yana! I know you know how to solve linear equations in one variables!!

OpenStudy (nincompoop):

yes, you should because that is part of what it will take to comprehend physical sciences concepts

OpenStudy (abhisar):

29 = 33 + a x 2.5 Find a in the above equation.

OpenStudy (nincompoop):

what makes this different from your algebra lessons is that it tends to be more application and analytical than learning the mechanics of solving x or y

OpenStudy (yanasidlinskiy):

Idk

OpenStudy (nincompoop):

yana, how well did you do in your algebra lessons?

OpenStudy (aaronandyson):

V = U + at Where , V =final velocity = 29m/s U = initial velocity = 33m/s a= accelerate = ? t = time = 2.5s 29 = 33 + a(2.5) 29 = 33 + 2.5a 29-33 = 2.5a -4 = 2.5a -4/2.5 = a -1.6 m/s^2 =a Since acceleration is negative we call this retardation. Therefore , retardation = -1.6 m/s^2 Now, To find the displacement , \[s = ut + \frac{ 1 }{ 2 }at\] Substitute the values and solve the equation and you'll get the distance traveled during this time.